OpenStudy (anonymous):

1. A consumer researcher wanted to know if customers really are influenced to buy more from sales clerks who smile. To test this, the clerks at eight stores in a large Canadian clothing chain were given special instructions at the start of a week and then sales over the week were recorded. Four of the stores were randomly selected to have the clerks receive instructions to be especially courteous and to smile a lot. Clerks at four other stores were simply instructed to be especially courteous. Sales (in thousands of dollars) for the four stores in the smile condition were 36, 40, 36, 44;

OpenStudy (anonymous):

Continued: sales for the other four stores in the control condition were 40, 31, 27, and 30. Do these results suggest that customers might buy more if they encounter smiling sales clerks? (Use the .05 level.)

OpenStudy (anonymous):

First you have to write a null hypothesis and an alternative hypothesis in this case Ho: clerks' sales = friendly clerks' sales Ha: clerks' sales < friendly clerks' sales

OpenStudy (anonymous):

So you find the conditions and assumptions and we are going to do a paired t -test

OpenStudy (anonymous):

We are going to find the difference in your numbers in the order you gave which would be 4,9,1,and 14

OpenStudy (anonymous):

Then you have to find the SE which is (stdDev/ the square root of n) = 5.716/ the square root of 4 = 5.716/2= 2.858

OpenStudy (anonymous):

After you would us the formula T (sub n-1) = (d-/\o)/SE * d= mean; /\o is almost always zero T(sub 3) = (7-0)/2.858)=4.142 Then enter into the calculator the option tcdf(4.142,100,3) and that should be your p-value. I think you may know how to do the rest...

OpenStudy (anonymous):

Hopefully I was help, this was the only formula I could work with the info given because all the rest involve different information such as the mean and stdDev and probability given. I assumed a few things and because of that you may have to use your own judgement or even ask your teacher to check the answer